Find the exact solution to the equation below.
52e^5x+3=26

Answer by stanbon(75887)

Find the exact solution to the equation below.
52e^(5x+3)=26
-------
e^(5x+3) = 1/2
-----
5x+3 = ln(1/2)
---
5x+3 = -0.6931
5x = -3.6931
x = -0.7386
-----
Cheers,
Stan H.
------------